# Disruptions of stars and binary systems on chaotic orbits in the axisymmetric Milky Way center

### v1

@@ Abstract
pubyear: 2024pagerange: Disruptions of stars and binary systems on chaotic orbits in the axisymmetric Milky Way center20
@@ 1 Introduction
Disrupted stars are observed as Tidal Disruption Events (TDEs), for which the bound debris forms a tightly bound accretion disk around the MBH (Hills, 1975; Rees, 1988). The formation and accretion of this disk causes a bright transient flare, observable in distant galaxies (van Velzen et al., 2020; Gezari, 2021), which fades over the course of months to years. Early efforts (Frank & Rees, 1976) to understand TDE rates led to the formulation of the stellar loss cone theory, which quantifies (Lightman & Shapiro, 1977; Cohn & Kulsrud, 1978) the collisional rate of disruption of stars which are are scattered (by two-body encounters) onto those radial orbits vulnerable to disruption.
@@
We can compare this collisionless channel for disruptions, to the much more widely studied collisional mechanism of scattering systems via a random walk of 2-body encounters (Frank & Rees, 1976; Lightman & Shapiro, 1977; Merritt, 2013b). The two channels produce not just different rates, but also different properties of the disruption and observational consequences. Collisionless loss cone refilling has been previously studied in Magorrian & Tremaine (1999); Merritt & Poon (2004); Khan et al. (2013); Vasiliev & Merritt (2013); Vasiliev et al. (2015) and Kaur & Stone (2024), and is sometime visualized as a loss-wedge - an expanded version of the usual loss-cone that has grown to encompass a much larger part of angular momentum space. We expand on the method and scope of these previous works, highlighting some specific changes in methodology that better capture the true physics and subsequent rates. We do this, for the first time, in the context of a detailed and up to date model of the Milky Way galactic center drawn from Sormani et al. (2022).
@@ 2.1 Variation in angular momentum
$\bm{\dot{l}}=\dot{l_{\theta}}\bm{\hat{\phi}}+l_{\theta}\bm{\dot{\hat{\phi}}}+\frac{\mathrm{c}_{\theta}l_{z}}{\mathrm{s}_{\theta}^{2}}\dot{\theta}\bm{\hat{\theta}}-\frac{l_{z}}{\mathrm{s}_{\theta}}\bm{\dot{\hat{\theta}}}=\left(\dot{l_{\theta}}-\frac{\mathrm{c}_{\theta}}{\mathrm{s}_{\theta}}\dot{\phi}l_{z}\right)\bm{\hat{\phi}}-\frac{1}{\mathrm{s}_{\theta}}\dot{l_{z}}\bm{\hat{\theta}}$ (12)
@@ 2.3 Close encounters with a central massive object
In a general potential there is no analytic expression for the apoapse or periapse radii, though they can be solved numerically by finding $r$ that satisfies equation 7 when $\dot{r}=0$. However, if the innermost part of the potential is dominated by a point-like MBH of mass $M$, then the local potential is approximately Keplerian [^3] At very small distances, a factor of a few Schwarzchild radii, relativistic effects can become important – and similarly for speeds approaching a significant fraction of the speed of light. Such conditions are not modeled here.. This small region is often denoted the sphere of influence, with $r\ll r_{\rm inf}$, the radius of influence.
@@ 3 Model of the Galactic center
We use a static density-potential pair, such that any orbit in our potential conserves energy. For much of this work, we employ the fiducial model of Sormani et al. (2022, S22 hereafter), which uses a NSC model from Chatzopoulos et al. (2015) and a NSD model fitted to the actions of stars observed in the GC. Following S22 we use the AGAMA code (Vasiliev, 2019), which has many capabilities that will be crucial for this work [^4] AGAMA includes a publicly accessible reproduction of the computation of the self-consistent NSD potential and DF performed in S22 - https://github.com/GalacticDynamics-Oxford/Agama/blob/master/py/example_mw_nsd.py. Firstly, AGAMA allows us to iteratively converge a DF and potential-density pair to reach self-consistency.
@@
We also define a fixed NSC DF from an updated model for the MW NSC (Vasiliev et al. in prep) - which is important for sampling orbital properties of systems in the NSC later. We do not perform the self-consistency with regard to this DF and thus it does not affect the GC potential. This means this DF and our GC model they are likely to be mildly inconsistent. However this DF is fitted to NSC data and we trust that it will provide a good approximation to the real kinematics in the NSC.
@@
We also show in Figure 2 the physical scales of our GC model, and how they vary with $R_{\rm c}$. The first panel shows the conversion between energy, $\epsilon$, and $R_{\rm c}$. We also show $\Phi_{0}$, the central potential excluding the MBH. Where the two lines meet roughly defines the influence radius and where the MBH potential comes to dominate on sub-parsec scales. The mass profile similarly shows the MBH (with mass $4\cdot 10^{6}M_{\odot}$) dominates the inner parsec, but is quickly superseded by the NSC (a few $10^{7}M_{\odot}$) up to about 10 parsec, the the NSD ($10^{9}M_{\odot}$) out to hundreds of parsecs, beyond which the bulge dominates. The characteristic timescale of orbits, $T_{\rm c}$, is the period of a circular orbit. $T_{\rm c}$ is a little under 100 kyr at the radius of influence and increases rapidly moving outwards ($\underset{\sim}{\propto}R_{\rm c}$). The charecteristic (and maximal) angular momentum of a circular orbit, $l_{\rm c}$, also increases sharply with radius (also $\underset{\sim}{\propto}R_{\rm c}$).
@@ 4 Trajectories
Figure 3: Examples of 5 closed trajectories in the GC potential, all with $R_{\rm c}(\epsilon)=50$ pc and $l_{z}=0$. For each we show the orbit, in units of $R_{\rm c}$, in the $x,\ z$ plane (upper left) and $R,\ z$ plane (lower left). The larger right-hand panel shows the orbit in $\theta$ (normalized to run from 1 at the north pole, at the 0 equator and -1 at the south pole) and $l_{\theta}$, normalized by the maximum angular momenta for a given energy $l_{\rm c}(\epsilon)$. Apoapsides are denoted by circles, and periapsides by diamonds, and numbered sequentially with 0 being the initial apoapse.
@@
As $l\geq l_{z}$ always, any trajectory that can experience disruption must have $l_{z}<l_{\rm t}$. Because tidal radii are small compared to galactic scales (for example $r_{\rm t}$ is typically $\lesssim$ AU for a hypervelocity star progenitor, or $\sim R_{\odot}$ for a TDE), $l_{\rm t}$ is tiny compared to typical angular momenta of most trajectories. However, while the average angular momentum of a trajectory may be large, for diving orbits $l_{\theta}$ can vary over orders of magnitude and may approach values $<l_{\rm t}$ for some orbits.
@@
We have two conserved quantities, $\epsilon$ and $l_{z}$, and a coordinate which contains no information, $\phi$, and thus the 6D phase space reduces to 3D. Many authors take these dimensions to be $R,\ z$ and $v_{R}$ (the velocity in the $R$ direction) [^5] Using conservation of energy, we could substitute any of these out for $v_{z}$, the vertical velocity.. Here we instead employ a different set of coordinates, inspired in part by Magorrian & Tremaine (1999), examining the polar angle $\theta$, and its corresponding angular momentum $l_{\theta}$ at apo- or peri-apse (where $\dot{r}=0$ and $\ddot{r}$ is negative or positive respectively). Either set of coordinates is sufficient to fully define the state (except for an arbitrary offset in $\phi$).
@@ 4.1 Surfaces of Section (SoS)
In Figure 3, we show 5 example trajectories in our GC potential, all with $R_{\rm c}(\epsilon)=50$ pc [^6] If we compare families of trajectories with the same specific energy $\epsilon$ then $R_{\rm c}(\epsilon)$ and thus $l_{\rm c}(\epsilon)$ are fixed for all trajectories in this family. We generally use $R_{\rm c}$ as a more intuitively accessible proxy for $\epsilon$ in our plots and discussion.. We show only examples with $l_{z}=0$, which prevents the trajectory from rotating around the z-axis. Hence we can set $y=0$ and $\dot{y}=0$ at all times, allowing easier visualization. For $l_{z}\neq 0$ trajectories, the $(x,z)$ plane becomes a less useful visualization, but the $(R,z)$ plane retains its clarity.
@@
Another distinct way to separate trajectories is by whether they always have a definite or reversible sense of polar rotation near apoapse. Some trajectories are capable of passing through $l_{\theta}=0$ at or near apoapse, corresponding a change of sign of $\dot{\theta}$ and a doubling back of the trajectory, which we would say have reversible polar rotation. Other trajectories maintain non-zero $l_{\theta}$ and $\dot{\theta}$ near apoapse, and thus cannot double back on themselves and continue with the same definite sense of polar rotation [^7] There is a sub-class of definite trajectories, that are called centrophillic because they pass directly through the origin. They have many interesting properties, including mimicking the orbital frequency harmonics of the reversible trajectories. However as our potential is singular at the origin they are not stable and do not appear in our analyses..
Regular closed trajectories (excluding higher harmonics) can be described by their sense of rotation and the number of apoapsides - which we can denote as $\mathcal{D}_{n}$ and $\mathcal{R}_{n}$ respectively. In Figure 3 we show, in order, the first two closed definite trajectories and the first three closed reversible trajectories: $\mathcal{D}_{2}$, (also commonly termed a box or loop trajectory), $\mathcal{D}_{3}$, $\mathcal{R}_{2}$ (or banana trajectory), $\mathcal{R}_{3}$ (or fish trajectory), and $\mathcal{R}_{4}$ (or pretzel trajectory). Definite trajectories require $n$ orbits to close, whilst reversible trajectories need $2(n-1)$ orbits.
@@
Not all closed trajectories (and associated islands of regular trajectories) exist for a given potential and integrals of motion. In particular, for a potential with a central singularity, such as ours, trajectories that would come near the origin are likely to diverge and become chaotic [^8] The exception to this is orbits that remain deep within the sphere of influence (i.e. $R_{\rm c}\ll r_{\rm infl}$) which experience an essentially Keplerian potential and are thus regular..
@@
Figure 4: Examples of 9 trajectories integrated in our GC potential with $R_{c}=50$ pc and $l_{z}=0$. We integrate each over $\sim$100 orbits, then perturb the initial conditions slightly and integrate over $\sim$1000 orbits. In the 9 small panels at the top we show each trajectory in $R$ and $z$ space (in units of $R_{\rm c}$) with the unperturbed trajectory shown as the darker foreground line and the perturbed trajectories shown in the lighter background region. We also show the $l_{\theta},\ \theta$ SoS at apoapse (middle panel) and periapse (bottom panel), for all integrated trajectories. The large numbered points in the apoapse SoS show the initial state of each unperturbed trajectory. The colours are consistent for each trajectory, and are also set by the initial state.
@@ 4.2 Classifying chaotic orbits
Figure 5: The full range of possible orbits with $R_{\rm c}$ of 50 pc and $l_{z}=0$, shown in the reduced space of $|l_{\theta}/l_{\rm c}|$ and $|\theta-\frac{\pi}{2}|/\frac{\pi}{2}$ (which is 0 at the equator and 1 at the poles). We show 2,000 random trajectories integrated over approximately 300 orbits. The leftmost column shows the state at apo- and peri- apse, coloured by the initial $l_{\theta,\rm p}$ and $\theta_{\rm p}$, with chaotic orbits ($\lambda>0.01$) in black. The second column shows the periapse space coloured by various properties of the orbit. $n_{\rm visit}$ is the number of cells, in a 30 by 30 grid over the space shown, visited by the trajectory - small values correspond to regular orbits that repeat their motion. $r_{\rm min}$ is the smallest radius visited by the orbit, here in units of $R_{\rm c}$. We also show the smallest (magnitude) $|l_{\theta}|_{\rm min}$ that the trajectory attains, and the value of the lyapunov exponent (in units of $T_{\rm c}^{-1}$) for chaotic orbits. In the righthand columns we show, in order, the 8 trajectories in $R$ and $z$ space (in units of $R_{\rm c}$) which visit the smallest number of cells - showing the regular orbits (and in some cases their harmonics) which are correspondingly numbered in the top left panel. Of these 5 are closed trajectories (0 - $\mathcal{D}_{2}$, 1 - $\mathcal{R}_{3}$, 2 - $\mathcal{R}_{2}$, 4 - $\mathcal{R}_{4}$ and 5 - $\mathcal{D}_{3}$) whilst the others are higher harmonics (3 & 6 - of $\mathcal{R}_{2}$ and 7 - of $\mathcal{R}_{3}$).

### v3

# Disruptions of stars and binary systems on chaotic orbits in an axisymmetric Milky Way center

@@ Abstract
pubyear: 2024pagerange: Disruptions of stars and binary systems on chaotic orbits in an axisymmetric Milky Way center20
@@ 1 Introduction
Disrupted stars are observed as Tidal Disruption Events (TDEs), for which the bound debris forms a tightly bound accretion disk around the MBH (Hills, 1975; Rees, 1988). The formation and accretion of this disk causes a bright transient flare, observable in distant galaxies (van Velzen et al., 2020; Gezari, 2021), which fades over the course of months to years. Early efforts (Frank & Rees, 1976) to understand TDE rates led to the formulation of the stellar loss-cone theory, which quantifies (Lightman & Shapiro, 1977; Cohn & Kulsrud, 1978) the collisional rate of disruption of stars which are are scattered (by two-body encounters) onto those radial orbits vulnerable to disruption.
@@
We can compare this collisionless channel for disruptions, to the much more widely studied collisional mechanism of scattering systems via a random walk of 2-body encounters (Frank & Rees, 1976; Lightman & Shapiro, 1977; Merritt, 2013b). The two channels produce not just different rates, but also different properties of the disruption and observational consequences. Collisionless loss-cone refilling has been previously studied in Magorrian & Tremaine (1999); Merritt & Poon (2004); Khan et al. (2013); Vasiliev & Merritt (2013); Vasiliev et al. (2015) and Kaur & Stone (2024), and is sometime visualized as a loss wedge - an expanded version of the usual loss cone that has grown to encompass a much larger part of angular momentum space. We expand on the method and scope of these previous works, highlighting some specific changes in methodology that better capture the true physics and subsequent rates. We do this, for the first time, in the context of a detailed and up to date model of the Milky Way galactic center drawn from Sormani et al. (2022).
@@ 2.1 Variation in angular momentum
$\bm{\dot{l}}=\dot{l}_{\theta}\bm{\hat{\phi}}+l_{\theta}\bm{\dot{\hat{\phi}}}+\left(\frac{\mathrm{c}_{\theta}l_{z}}{\mathrm{s}_{\theta}^{2}}\dot{\theta}-\frac{\dot{l}_{z}}{\mathrm{s}_{\theta}}\right)\bm{\hat{\theta}}-\frac{l_{z}}{\mathrm{s}_{\theta}}\bm{\dot{\hat{\theta}}}=\left(\dot{l}_{\theta}-\frac{\mathrm{c}_{\theta}}{\mathrm{s}_{\theta}}\dot{\phi}l_{z}\right)\bm{\hat{\phi}}-\frac{\dot{l}_{z}}{\mathrm{s}_{\theta}}\bm{\hat{\theta}}$ (12)
@@ 2.3 Close encounters with a central massive object
In a general potential there is no analytic expression for the apoapse or periapse radii, though they can be solved numerically by finding $r$ that satisfies equation 7 when $\dot{r}=0$. However, if the innermost part of the potential is dominated by a point-like MBH of mass $M$, then the local potential is approximately Keplerian. [^3] At very small distances, a factor of a few Schwarzchild radii, relativistic effects can become important – and similarly for speeds approaching a significant fraction of the speed of light. Such conditions are not modeled here. This small region is often denoted the sphere of influence, with $r\ll r_{\rm inf}$, the radius of influence.
@@ 3 Model of the Galactic center
We use a static density-potential pair, such that any orbit in our potential conserves energy. For much of this work, we employ the fiducial model of Sormani et al. (2022, S22 hereafter), which uses a NSC model from Chatzopoulos et al. (2015) and a NSD model fitted to the actions of stars observed in the GC. Following S22 we use the AGAMA code (Vasiliev, 2019), which has many capabilities that will be crucial for this work. [^4] AGAMA includes a publicly accessible reproduction of the computation of the self-consistent NSD potential and DF performed in S22 - https://github.com/GalacticDynamics-Oxford/Agama/blob/master/py/example_mw_nsd.py Firstly, AGAMA allows us to iteratively converge a DF and potential-density pair to reach self-consistency.
@@
We also define a fixed NSC DF from an updated model for the MW NSC (Vasiliev et al. in prep) which follows a similar method to S22 fitting NSC stars in action space. As discussed in Appendix A this produces a distribution function that is centrally tangentially biased and which we will use to sample the orbital properties of systems in the NSC later. We do not perform the same iterative updating of the NSC potential to reach a self-consistent solution. Therefore, this NSC DF and the NSC potential from S22 that is used in our GC model are likely to be mildly inconsistent. However, as this DF is fitted to NSC data, we use it as a best available approximation to the real kinematics in the NSC.
@@
We also show in Figure 2 the physical scales of our GC model, and how they vary with $R_{\rm c}$. The first panel shows the conversion between energy, $\epsilon$, and $R_{\rm c}$. We also show $\Phi_{0}$, the central potential excluding the MBH. Where the two lines meet roughly defines the influence radius and where the MBH potential comes to dominate on sub-parsec scales. The mass profile similarly shows the MBH (with mass $4\cdot 10^{6}M_{\odot}$) dominates the inner parsec, but is quickly superseded by the NSC (a few $10^{7}M_{\odot}$) up to about 10 parsec, the NSD ($10^{9}M_{\odot}$) out to hundreds of parsecs, beyond which the bulge dominates. The characteristic timescale of orbits, $T_{\rm c}$, is the period of a circular orbit. $T_{\rm c}$ is a little under 100 kyr at the radius of influence and increases rapidly moving outwards ($\underset{\sim}{\propto}R_{\rm c}$). The charecteristic (and maximal) angular momentum of a circular orbit, $l_{\rm c}$, also increases sharply with radius (also $\underset{\sim}{\propto}R_{\rm c}$).
@@ 4 Trajectories
Figure 3: Examples of 5 closed trajectories in the GC potential, all with $R_{\rm c}(\epsilon)=50$ pc and $l_{z}=0$. We label them by their classification, $\mathcal{D}_{\rm n}$ or $\mathcal{R}_{\rm n}$ (see section 4.1). For each we show the trajectory, in units of $R_{\rm c}$, in the $x,\ z$ plane (upper left) and $R,\ z$ plane (lower left). The larger right-hand panel shows the evolution in $\theta$ (normalized to run from 1 at the north pole, at the 0 equator and -1 at the south pole) and $l_{\theta}$ (normalized by the maximum angular momenta for a given energy $l_{\rm c}(\epsilon)$). Apoapsides are denoted by circles, and periapsides by diamonds, and numbered sequentially with 0 being the initial apoapse. For reversible ($\mathcal{R}_{\rm n}$) trajectories we only show the first half of a full closed trajectory, as after this they double back on themselves in physical space and they mirror (through $l_{\theta}=0$) their path in $(\theta,l_{\theta})$ space.
@@
As $l\geq l_{z}$ always, any trajectory that can experience disruption must have $l_{z}<l_{\rm t}$. Because tidal radii are small compared to galactic scales (for example $r_{\rm t}$ is typically $\lesssim$ 100 AU for a hypervelocity star progenitor, or 1 AU for a TDE), $l_{\rm t}$ is tiny compared to typical angular momenta of most trajectories. However, while the average angular momentum of a trajectory may be large, for diving orbits $l_{\theta}$ can vary over orders of magnitude and may approach values $<l_{\rm t}$ for some orbits.
@@
We have two conserved quantities, $\epsilon$ and $l_{z}$, and a coordinate which contains no information, $\phi$, and thus the 6D phase space reduces to 3D. Many authors take these dimensions to be $R,\ z$ and $v_{R}$ (the velocity in the $R$ direction). [^5] Using conservation of energy, we could substitute any of these out for $v_{z}$, the vertical velocity. Here we instead employ a different set of coordinates, inspired in part by Magorrian & Tremaine (1999), examining the polar angle $\theta$, and its corresponding angular momentum $l_{\theta}$ at apo- or peri-apse (where $\dot{r}=0$ and $\ddot{r}$ is negative or positive respectively). Either set of coordinates is sufficient to fully define the state (except for an arbitrary offset in $\phi$).
@@ 4.1 Surfaces of Section (SoS)
In Figure 3, we show 5 example trajectories in our GC potential, all with $R_{\rm c}(\epsilon)=50$ pc. [^6] If we compare families of trajectories with the same specific energy $\epsilon$ then $R_{\rm c}(\epsilon)$ and thus $l_{\rm c}(\epsilon)$ are fixed for all trajectories in this family. We generally use $R_{\rm c}$ as a more intuitively accessible proxy for $\epsilon$ in our plots and discussion. We show only examples with $l_{z}=0$, which prevents the trajectory from rotating around the z-axis. Hence we can set $y=0$ and $\dot{y}=0$ at all times, allowing easier visualization. For $l_{z}\neq 0$ trajectories, the $(x,z)$ plane becomes a less useful visualization, but the $(R,z)$ plane retains its clarity.
@@
Another distinct way to separate trajectories is by whether they always have a definite or reversible sense of polar rotation near apoapse. Some trajectories are capable of passing through $l_{\theta}=0$ at or near apoapse, corresponding a change of sign of $\dot{\theta}$ and a doubling back of the trajectory, which we would say have reversible polar rotation. Other trajectories maintain non-zero $l_{\theta}$ and $\dot{\theta}$ near apoapse, and thus cannot double back on themselves and continue with the same definite sense of polar rotation. [^7] There is a sub-class of definite trajectories, that are called centrophillic because they pass directly through the origin. They have many interesting properties, including mimicking the orbital frequency harmonics of the reversible trajectories. However as our potential is singular at the origin they are not stable and do not appear in our analyses.
Regular closed trajectories (excluding higher harmonics) can be described by their sense of rotation and the number of apoapsides. Definite (reversible) trajectories with n apoapsides are denoted as $\mathcal{D}_{n}$ ($\mathcal{R}_{n}$). In Figure 3 we show, in order, the first two closed definite trajectories and the first three closed reversible trajectories: $\mathcal{D}_{2}$, (also commonly termed a box or loop trajectory), $\mathcal{D}_{3}$, $\mathcal{R}_{2}$ (or banana trajectory), $\mathcal{R}_{3}$ (or fish trajectory), and $\mathcal{R}_{4}$ (or pretzel trajectory). Definite trajectories require $n$ orbits to close, whilst reversible trajectories need $2(n-1)$ orbits.
@@
Not all closed trajectories (and associated islands of regular trajectories) exist for a given potential and integrals of motion. In particular, for a potential with a central singularity, such as ours, trajectories that would come near the origin are likely to diverge and become chaotic. [^8] The exception to this is orbits that remain deep within the sphere of influence (i.e. $R_{\rm c}\ll r_{\rm infl}$) which experience an essentially Keplerian potential and are thus regular.
@@
Figure 4: Examples of 9 trajectories integrated in our GC potential with $R_{c}=50$ pc and $l_{z}=0$. We integrate each over $\sim$100 orbits (solid line, large dots), then perturb the initial conditions slightly and integrate over $\sim$1000 orbits (translucent line, small dots). In the 9 small panels at the top we show each trajectory in $R$ and $z$ space (in units of $R_{\rm c}$). We also show the $(l_{\theta},\theta)$ SoS at apoapse (middle panel) and periapse (bottom panel), for all integrated trajectories. The large numbered points in the apoapse SoS show the initial state of each unperturbed trajectory.
@@ 4.2 Classifying chaotic orbits
Figure 5: The full range of possible orbits with $R_{\rm c}$ of 50 pc and $l_{z}=0$, shown in the reduced space of $|l_{\theta}/l_{\rm c}|$ and $|\theta-\frac{\pi}{2}|/\frac{\pi}{2}$ (which is 0 at the equator and 1 at the poles). We show 2,000 random trajectories integrated over approximately 300 orbits. The leftmost column shows the state at apo- and peri- apse, coloured by the initial $l_{\theta,\rm p}$ and $\theta_{\rm p}$, with chaotic orbits ($\lambda>0.01$) in black. The second column shows the periapse space coloured by various properties of the orbit. $n_{\rm visit}$ is the number of cells, in a 30 by 30 grid over the space shown, visited by the trajectory - small values correspond to regular orbits that repeat their motion. $r_{\rm min}$ is the smallest radius visited by the orbit, here in units of $R_{\rm c}$. We also show the smallest (magnitude) $|l_{\theta}|_{\rm min}$ that the trajectory attains, and the value of the lyapunov exponent, $\lambda$, (in units of $T_{\rm c}^{-1}$) for chaotic orbits. In the righthand columns we show, in order, the 8 trajectories in $R$ and $z$ space (in units of $R_{\rm c}$) which visit the smallest number of cells - showing the regular orbits (and in some cases their harmonics) which are correspondingly numbered in the top left panel. Of these 5 are closed trajectories (0 - $\mathcal{D}_{2}$, 1 - $\mathcal{R}_{3}$, 2 - $\mathcal{R}_{2}$, 4 - $\mathcal{R}_{4}$ and 5 - $\mathcal{D}_{3}$) whilst the others are higher harmonics (3 & 6 - of $\mathcal{R}_{2}$ and 7 - of $\mathcal{R}_{3}$).